OpenStudy (anonymous):

Solve 4^(x + 7) = 6^(x – 1). Write your answer in decimal form.

OpenStudy (kira_yamato):

4^(x + 7) = 6^(x – 1) (x + 7)ln4 = (x - 1) ln 6 xln4 + 7ln4 = xln6 - ln6 xln6 - xln4 = ln6 + 7ln4 x (ln6 - ln4) = ln6 + 7ln4 x = (ln6 + 7ln4) / (ln6 - ln4) Press this into your calculator for the answer

OpenStudy (anonymous):

Hey @Kira_Yamato are you sure that's the correct answer?

OpenStudy (anonymous):

Sorry, it's correct.